Find 4 consecutive even integers where the product of the two smaller numbers is 72 less than the product of the two larger numbers?
Answer by RAY100(1637)

Let 1st even integer = x, next (x+2) , next (x+4) , next ( x+6)
.
expression ,,,,,,(x)(x+2) = (x+4)(x+6) -72
.
x^2 +2x = x^2 +10x +24 -72
.
72 -24 = 10x -2x
.
48 = 8x
.
x=6
.
set is ,,,,,6,,,,,8,,,,10,,,12
.
checking
.
(6*8) = ( 10*12) -72
.
48 = 120 -72 = 48,,,,,,ok